a. Fill in the missing numbers in the following income statement: (Do not round intermediate calculations and round your answcan you please help me with this finance question
Answer #1
Sales 586900
Costs 364300
Depreciation 114100
EBIT 108500
Taxes(108500*22%) $23870
Net income $84630

OCF=Net income+Depreciation

=(84630+114100)=$198730

Depreciation tax shield=Depreciation*Tax rate

=114100*22%=$25102